OpenStudy (anonymous):

Find the minimum of f(x)=xsqrt(x+1) on the interval (-1,1)

OpenStudy (anonymous):

I got (uv)'=u'v+uv' u=x u'=1 v=sqrt(x+1) v'=1/2(x+1)^(-1/2)

OpenStudy (anonymous):

with that i got x(1/2)(x+1)^(-1/2)+sqrt(x+1) now idk what to do??

OpenStudy (solomonzelman):

you got the v' and u' right.

OpenStudy (anonymous):

yes, u'=1 and v'=(1/2)(x+1)^(-1/2)

OpenStudy (solomonzelman):

and the (uv)' together is correct also.

OpenStudy (anonymous):

\[v'=\frac{ 1 }{ 2 }(x+1)^{\frac{ -1 }{ 2 }}\]

OpenStudy (anonymous):

so i set it to 0?

OpenStudy (solomonzelman):

yes, you set (uv)' equal to zero.

OpenStudy (anonymous):

\[x \frac{ 1 }{ 2 }(x+1)^{\frac{ -1 }{ 2 }}+\sqrt{x+1}=0\]

OpenStudy (solomonzelman):

\(\large\color{black}{ 0=\frac{x}{2\sqrt{x+1}} +\sqrt{x+1} }\)

OpenStudy (solomonzelman):

I would play with the (uv)' for a bit though.

OpenStudy (anonymous):

should i devide both sides by \[\sqrt{x+1}\]?

OpenStudy (anonymous):

okay

OpenStudy (solomonzelman):

\(\Large\color{black}{ 0=\frac{x}{2\sqrt{x+1}} +\frac{2(x+1)}{2\sqrt{x+1}} }\) \(\Large\color{black}{ 0=\frac{x}{2\sqrt{x+1}} +\frac{2x+1}{2\sqrt{x+1}} }\) \(\Large\color{black}{ 0=\frac{3x+1}{2\sqrt{x+1}} }\)

OpenStudy (solomonzelman):

See what I did?

OpenStudy (solomonzelman):

now, multiply both sides times \(\Large\color{black}{ 2\sqrt{x+1} }\) .

OpenStudy (solomonzelman):

and, ... Are you sure that the interval is (-1,1) and NOT [-1,1] ?

OpenStudy (anonymous):

on the hw sheet it has -11,1 in () as (-1,1)

OpenStudy (anonymous):

but I am following what you have done so far

OpenStudy (solomonzelman):

So your interval is \(\Large\color{black}{ (-11,1) }\) ?

OpenStudy (anonymous):

and im left with 3x+1)=0

OpenStudy (solomonzelman):

yes, with 3x+1=0, so x equals?

OpenStudy (solomonzelman):

u there ?

OpenStudy (anonymous):

1/3=x

OpenStudy (solomonzelman):

-1/3, no?

OpenStudy (anonymous):

yes sorry, scrolled up for a second oh yes -1/3

OpenStudy (solomonzelman):

and again, your interval is (-11,1) correct?

OpenStudy (anonymous):

(-1,1) not -11

OpenStudy (solomonzelman):

Ohh. if it is \(\Large\color{black}{ (\color{red}{-1},1) }\), then -1, is not a critical number, even though it is in the domian of f(x), and f'(x) is undefined at it.

OpenStudy (solomonzelman):

Then, -1/3 would be your only critical number.

OpenStudy (anonymous):

would finding f'' get me the minimum then?

OpenStudy (solomonzelman):

No, you needed the critical numbers. And the smallest f ' (critical number) you get, that is the the absolute minimum.

OpenStudy (solomonzelman):

f '' (x) has to do with testing concavity, and it is different.

OpenStudy (anonymous):

okay! So than -1/3 is the only answer I need than. You are awesome thank you soo much!!!!

OpenStudy (solomonzelman):

You are not done.

OpenStudy (solomonzelman):

-1/3 is the critical number with which you would get an absolute minimum of the function, but the absolute minimum itself, is f(-1/3).

OpenStudy (solomonzelman):

Calculate f(-1/3)

OpenStudy (anonymous):

okay!!! f(-1/3)=1/3sqrt(4/3)

OpenStudy (anonymous):

\[f(\frac{ -1 }{ 3 })=(\frac{ 1 }{ 3 })\sqrt{\frac{4 }{ 3 }}\]

OpenStudy (anonymous):

can that be simplified?

OpenStudy (solomonzelman):

it is negative 1/3, not positive 1/3.

OpenStudy (anonymous):

oh yes

OpenStudy (anonymous):

it can \[f(\frac{ -1 }{ 3 })=(-\frac{ \sqrt{\frac{ 4 }{ 3 }} }{ 3 })\]

OpenStudy (anonymous):

ohhh i see it is negative than that means the 4/3 in sqrt is 2/3

OpenStudy (solomonzelman):